OpenStudy (anonymous):

why the heck am i getting this wrong? The electric field strength is 2.30×104N/C inside a parallel-plate capacitor with a 1.50mm spacing. An electron is released from rest at the negative plate.What is the electron's speed when it reaches the positive plate?

OpenStudy (anonymous):

Felectron=1.6*10^-19 * 2.3*10^4=3.68*10^-15 a=3.68*10^-15/9.109*10^-31=4.04*10^15 v=sqrt(2*4.04*10^15*1.5*10^-6)=110090 m/s mastering physics says it's wrong. i don't understand what i did wrong.

OpenStudy (shamim):

I think u can use as ur last equation v^2=u^2+2*a*s v^2=0^2+2*4.04*10^15*0.0015 v=?

OpenStudy (anonymous):

ok, thanks. usually my conversions get messed up for some reason.

OpenStudy (shamim):

U r most welcome

OpenStudy (shamim):

Anyway, glad to help u!!!